# Forget BIT, It is All about TOKEN: Towards Semantic Information Theory for LLMs

### v1

@@ IV-A Token-level Semantic Space
The token-level semantic space of a language is a probabilistic space $(\Omega,\mathscr{F},P)$, where $|\Omega|=N\geq 1$ is a set of all tokens, each of which is the atomic unit with specific semantics in this language, $\mathscr{F}\subseteq 2^{\Omega}$ is the $\sigma$-algebra, $P$ is the probability measure defined on $\mathscr{F}$.
@@ IV-B Token-level Semantic Vector Space
The token-level semantic vector space of a language is a probabilistic inner product space $\mathsf{S}=(\mathbb{S}^{N-1},\mathscr{F},\mu,\langle\cdot,\cdot\rangle)$, where $\mathbb{S}^{N-1}$ is a (N1)(N-1)-dimensional unit sphere, each $\mathbf{s}\in\mathbb{S}^{N-1}$ represents a semantic vector, $\mathscr{F}$ is a $\sigma$-algebra on $\mathbb{S}^{N-1}$, $\mu$ is a probability measure defined on $\mathscr{F}$, $\langle\cdot,\cdot\rangle$ is an inner product.
If we use $\mathbf{s_{1}}$ and $\mathbf{s}_{2}$ to denote two column vectors on $\mathbb{S}^{N-1}$, the inner product can be written as $\langle\mathbf{s}_{1},\mathbf{s}_{2}\rangle=\mathbf{s}_{1}^{T}\mathbf{s}_{2}$. The squared Euclidean distance is defined as $d_{e}^{2}(\mathbf{s}_{1},\mathbf{s}_{2})=\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}=(\mathbf{s}_{1}-\mathbf{s}_{2})^{T}(\mathbf{s}_{1}-\mathbf{s}_{2})$. The cosine similarity is defined as $\cos(\mathbf{s}_{1},\mathbf{s}_{2})=\mathbf{s}_{1}^{T}\mathbf{s}_{2}$. It is noticed that $\Omega$ in Definition 10 can only be mapped to $N$ points in $\mathbb{S}^{N-1}$. Let the set of semantic vector of tokens in $\mathcal{A}$ be $\mathcal{S}(\mathcal{A})\subset\mathbb{S}^{N-1}$ with $\forall\mathcal{A}\subseteq\Omega$. Thus, $\mu$ is an extension from $P$ such that $\mu(\mathcal{S}(\mathcal{A}))=P(\mathcal{A})$ if $\mathcal{A}\in\mathscr{F}_{\mathsf{S}}$, otherwise $\mu(\mathcal{S}(\mathcal{A}))=0$.
Many works suggest that the semantic vector space should be a more complex low dimensional manifold. In practice, however, the Euclidean distance and cosine similarity remain the most widely used metrics, because of its simplicity in computation and adequate performance. Therefore, we argue that defining the semantic vector space directly on $\mathbb{S}^{N-1}$ strikes an effective trade-off between accuracy and computational efficiency.
@@
$d_{s}^{2}(\mathsf{S},\mathsf{S}^{\prime})=\min_{\pi\in\Pi(\mu,\nu)}\int_{\mathsf{S}\times\mathsf{S}^{\prime}}\int_{\mathsf{S}\times\mathsf{S}^{\prime}}\left|\mathbf{s}_{1}^{T}\mathbf{s}^{\prime}_{1}-\mathbf{s}_{2}^{T}\mathbf{s}^{\prime}_{2}\right|^{2}d\pi(\mathbf{s}_{1},\mathbf{s}_{2})d\pi(\mathbf{s}^{\prime}_{1},\mathbf{s}^{\prime}_{2}),$ (34)
@@
The relationship between semantic space and semantic vector space is similar to the relationship between information theory and signal processing. Information theory, based on probability theory, is a framework for understanding the nature and limits of information compression, transmission, and storage. However, it is not particularly concerned with the specific methods of implementation in practice [16]. Signal processing, on the other hand, represents information as vectors in $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$, making it suitable for sensing, transmission, and storage in physical media. This representation enables a vast body of mathematical theory to be applied to the design of efficient algorithms for practical sensing, communication, and storage systems [82].
@@ IV-C Semantic Compression/De-dimensionality
In information theory, the objective of source coding is to use as few bits as possible to represent a source symbol, such that the source message can be exactly recovered for lossless compression or recovered within a given distortion for lossy compression [43]. According to Definition 11, however, $|\Omega|=N$ implies $\mathbb{S}^{N-1}$ is a very high dimension sphere such that the direct computation on $\mathsf{S}$ is still not practical. Extensive experimental results suggest that the choice of dimensionality for a semantic vector space involves a crucial trade-off, implying the existence of an optimal range or “sweet spot” [83]. In this case, the semantic compression is the compression of the entire semantic space, i.e., dimension reduction that preserves cosine similarity.
@@
Let $\epsilon\in(0,1)$ and $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{M}\}\in\mathbb{S}^{N-1}$, if $m\geq\frac{C}{\epsilon^{2}}\log M$, there exists a matrix $\mathbf{A}\in\mathbb{R}^{m\times N}$ such that:
$|\mathbf{s}_{i}^{T}\mathbf{s}_{j}-\mathbf{s}_{i}^{T}\mathbf{P}\mathbf{s}_{j}|\leq\epsilon,\quad\forall i,j\in\{1,\ldots,M\},$ (35)
@@
According to JL lemma, the dimensionality of the semantic vector space can be reduced from $N$ to $m\geq\frac{C}{\epsilon^{2}}\log M$. As aforementioned, each semantic vector can be seen as a real signal vector which should be very sparse in $\mathbb{S}^{N-1}$. Inspired by compressive sensing, the cosine similarity based JL lemma can be improved by applying restricted isometry property (RIP). Let $\mathbf{A}$ be a matrix satisfying $(k,\delta)$-RIP, that is
@@
for all $k$-sparse $\mathbf{s}\in\mathbb{S}^{N-1}$, i.e., $\|\mathbf{s}\|_{0}\leq k$. The following result is established in [86].
@@
Let $\eta,\epsilon\in(0,1)$, $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{M}\}\in\mathbb{S}^{N-1}$, and $\mathbf{A}\in\mathbb{R}^{m\times N}$ be $(k,\delta)$-RIP with $\delta\leq\epsilon/4$ and $k\geq 40\log\frac{4M}{\eta}$. Let $\bm{\sigma}$ a Rademacher sequence, i.e., uniformly distributed on $\{-1,1\}^{N}$. Then, with probability exceeding $1-\eta$,
$|\mathbf{s}_{i}^{T}\mathbf{s}_{j}-\mathbf{s}_{i}^{T}\mathbf{D}_{\bm{\sigma}}\mathbf{P}\mathbf{D}_{\bm{\sigma}}\mathbf{s}_{j}|\leq\epsilon,\quad\forall i,j\in\{1,\ldots,M\},$ (37)
@@
According to the theory of compressive sensing, the $m\times N$ partial Gaussian matrix can be used with
$m\geq\frac{C}{\epsilon^{2}}\log\frac{M}{\eta}\log N,$ (38)
@@
Recalling Definition 12, the distortion of semantic compression can be evaluated by the distance of two semantic vector spaces. Let $\mathsf{S}$ be the original semantic vector space on $\mathbb{S}^{N-1}$ and $\mathsf{S}^{\prime}$ on $\mathbb{S}^{m}$ with $1\leq m<N-1$, the distortion of semantic compression can be written as
$d_{s}^{2}(\mathsf{S},\mathsf{S}^{\prime})=\min_{\pi\in\Pi(\mu,\mu^{\prime})}\int_{\mathsf{S}\times\mathsf{S}^{\prime}}\int_{\mathsf{S}\times\mathsf{S}^{\prime}}\left|\mathbf{s}^{T}\mathbf{s}^{\prime}-\mathbf{s}^{T}\mathbf{P}\mathbf{s}^{\prime}\right|^{2}d\pi(\mathbf{s},\mathbf{As})d\pi(\mathbf{s}^{\prime},\mathbf{As}^{\prime}),$ (39)
where $\mathbf{A}$ is a $m\times N$ projection matrix and $\mathbf{P}=\mathbf{A}^{T}\mathbf{A}$. The following theorem can be established by applying Lemma 1 or Theorem 3 directly.
@@ IV-D Semantic Embedding/Vectorization for Next-token Prediction
In practice, we typically select a proper dimension $m$ to directly perform the semantic embedding or vectorization. In the following, we will discuss information-theoretically optimal approach. It is natural to understand that the semantics of an utterance highly depend on the speaker’s intended goal, i.e., the downstream task in machine learning. Therefore, for a token sequence with length $n$, the semantic embedding is a mapping $f:\Omega^{n}\to(\mathbb{S}^{m})^{n}$, such that a loss functional $L(f)$, defined by the downstream task, is minimized.
From the perspective of LLMs, the objective is to predict the next token based on the prompt and the parameterized memory. Therefore, $L(f)$ should be designed to best facilitate of achieving this goal. Let $X_{1:n}$ be a token sequence, $S_{1:n}$ be the corresponding semantic vector representation of $X_{1:n}$. For the task of the next token prediction, $S_{t}$ should contain all the information in $X_{1:t}$ which is useful to predict $X_{t+1:n}$. From the perspective of information theory, the optimal semantic encoder for next token prediction should be the solution of the following problem:
$\max_{S_{t}=f(X_{1:t})}I(X_{t+1:n};S_{t}|S_{1:t-1}),\quad 1\leq t\leq n\in\mathbb{N}.$ (40)
The condition means $S_{t}$ only contains new information for predicting $X_{t+1:n}$ which is not contained in $S_{1:t-1}$.
@@
$I^{*}(X_{n:1}\to S_{1:n})\leq\sum_{t=1}^{n}\max_{S_{t}=f(X_{1:t})}I(X_{t+1:n};S_{t})\leq\sum_{t=1}^{n}\sum_{k=1}^{n-t}\max_{S_{t}=f(X_{1:t})}I(X_{t+k};S_{t}).$ (42)
Inspired by the idea of predictive coding in information theory [87, 88], the CPC is proposed for semantic embedding in [48], which is also adopt in OpenAI [49]. Let $Z_{1:n}$ be the latent representation of $X_{1:n}$ with $Z_{t}=g_{\text{ENC}}(X_{t})$, $S_{1:n}$ be the semantic vector obtained by CPC, which is defined as $S_{t}=g_{\text{AR}}(Z_{1:t-1})$. The training process of CPC is to solve the following optimization problem:
$\sum_{k=1}^{n-t}\max_{S_{t}=f(X_{1:t})}I(X_{t+k};S_{t}).$ (43)
@@ V-A TV-VAR based AR-LLMs
Let $\mathbf{s}_{t}$ with $t=1,\ldots,n$ and $\mathbf{u}_{t}$ with $t=n+1,\ldots,T$ be sample vectors of random variables $S_{t}$ and $U_{t}$. To simplify the notation, we let $\mathbf{u}_{t}=\mathbf{s}_{t}$ for $t=1,\ldots,n$. We then have the following definition.
@@
$\mathbf{u}_{t}=\arg\mathrm{softmax}\left(\frac{1}{\Xi}\tilde{\mathbf{u}}_{1:N}^{T}\left(\sum_{j=1}^{t-1}\mathbf{A}_{tj}\mathbf{u}_{j}\right)\right),\quad t=n+1,\ldots,T,$ (44)
where $\mathbf{A}_{tj}$ is the coefficient matrix, $\tilde{\mathbf{u}}_{1:N}$ are all possible token vectors in $\mathcal{S}(\Omega)$, and $\Xi$ is the sampling temperature.
In contrast to the standard VAR model [50], $\mathbf{A}_{tj}$ is time-variant, which is very difficult to estimate in practice.

### v2

@@ IV-A Token-level Semantic Space
The token-level semantic space of a language is a probabilistic space $(\Omega,\mathscr{F},P)$, where $|\Omega|=M\geq 1$ is a set of all tokens, each of which is the atomic unit with specific semantics in this language, $\mathscr{F}\subseteq 2^{\Omega}$ is the $\sigma$-algebra, $P$ is the probability measure defined on $\mathscr{F}$.
@@ IV-B Token-level Semantic Vector Space
The token-level semantic vector space of a language is a probabilistic inner product space $\mathsf{S}=(\mathbb{S}^{M-1},\mathscr{F},\mu,\langle\cdot,\cdot\rangle)$, where $\mathbb{S}^{M-1}$ is a (M1)(M-1)-dimensional unit sphere, each $\mathbf{s}\in\mathbb{S}^{M-1}$ represents a semantic vector, $\mathscr{F}$ is a $\sigma$-algebra on $\mathbb{S}^{M-1}$, $\mu$ is a probability measure defined on $\mathscr{F}$, $\langle\cdot,\cdot\rangle$ is an inner product.
If we use $\mathbf{s_{1}}$ and $\mathbf{s}_{2}$ to denote two column vectors on $\mathbb{S}^{M-1}$, the inner product can be written as $\langle\mathbf{s}_{1},\mathbf{s}_{2}\rangle=\mathbf{s}_{1}^{T}\mathbf{s}_{2}$. The squared Euclidean distance is defined as $d_{e}^{2}(\mathbf{s}_{1},\mathbf{s}_{2})=\|\mathbf{s}_{1}-\mathbf{s}_{2}\|^{2}=(\mathbf{s}_{1}-\mathbf{s}_{2})^{T}(\mathbf{s}_{1}-\mathbf{s}_{2})$. The cosine similarity is defined as $\cos(\mathbf{s}_{1},\mathbf{s}_{2})=\mathbf{s}_{1}^{T}\mathbf{s}_{2}$. It is noticed that $\Omega$ in Definition 10 can only be mapped to $M$ points in $\mathbb{S}^{M-1}$. Let the set of semantic vector of tokens in $\mathcal{A}$ be $\mathcal{S}(\mathcal{A})\subset\mathbb{S}^{M-1}$ with $\forall\mathcal{A}\subseteq\Omega$. Thus, $\mu$ is an extension from $P$ such that $\mu(\mathcal{S}(\mathcal{A}))=P(\mathcal{A})$ if $\mathcal{A}\in\mathscr{F}_{\mathsf{S}}$, otherwise $\mu(\mathcal{S}(\mathcal{A}))=0$.
Many works suggest that the semantic vector space should be a more complex low dimensional manifold. In practice, however, the Euclidean distance and cosine similarity remain the most widely used metrics, because of its simplicity in computation and adequate performance. Therefore, we argue that defining the semantic vector space directly on $\mathbb{S}^{M-1}$ strikes an effective trade-off between accuracy and computational efficiency.
@@
$d_{s}^{2}(\mathsf{S},\mathsf{S}^{\prime})=\min_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{S}^{M-1}\times\mathbb{S}^{M^{\prime}-1}}\int_{\mathbb{S}^{M-1}\times\mathbb{S}^{M^{\prime}-1}}\left|\mathbf{s}_{1}^{T}\mathbf{s}^{\prime}_{1}-\mathbf{s}_{2}^{T}\mathbf{s}^{\prime}_{2}\right|^{2}d\pi(\mathbf{s}_{1},\mathbf{s}_{2})d\pi(\mathbf{s}^{\prime}_{1},\mathbf{s}^{\prime}_{2}),$ (34)
@@
The relationship between semantic space and semantic vector space is similar to the relationship between information theory and signal processing. Information theory, based on probability theory, is a framework for understanding the nature and limits of information compression, transmission, and storage. However, it is not particularly concerned with the specific methods of implementation in practice [16]. Signal processing, on the other hand, represents information as vectors in $\mathbb{R}^{M}$ or $\mathbb{C}^{M}$, making it suitable for sensing, transmission, and storage in physical media. This representation enables a vast body of mathematical theory to be applied to the design of efficient algorithms for practical sensing, communication, and storage systems [82].
@@ IV-C Semantic Compression/De-dimensionality
In information theory, the objective of source coding is to use as few bits as possible to represent a source symbol, such that the source message can be exactly recovered for lossless compression or recovered within a given distortion for lossy compression [43]. According to Definition 11, however, $|\Omega|=M$ implies $\mathbb{S}^{M-1}$ is a very high dimension sphere such that the direct computation on $\mathsf{S}$ is still not practical. Extensive experimental results suggest that the choice of dimensionality for a semantic vector space involves a crucial trade-off, implying the existence of an optimal range or “sweet spot” [83]. In this case, the semantic compression is the compression of the entire semantic space, i.e., dimension reduction that preserves cosine similarity.
@@
Let $\epsilon\in(0,1)$ and $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{K}\}\in\mathbb{S}^{M-1}$, if $m\geq\frac{C}{\epsilon^{2}}\log K$, there exists a matrix $\mathbf{A}\in\mathbb{R}^{m\times M}$ such that:
$|\mathbf{s}_{i}^{T}\mathbf{s}_{j}-\mathbf{s}_{i}^{T}\mathbf{P}\mathbf{s}_{j}|\leq\epsilon,\quad\forall i,j\in\{1,\ldots,K\},$ (35)
@@
According to JL lemma, the dimensionality of the semantic vector space can be reduced from $M$ to $m\geq\frac{C}{\epsilon^{2}}\log K$. As aforementioned, each semantic vector can be seen as a real signal vector which should be very sparse in $\mathbb{S}^{M-1}$. Inspired by compressive sensing, the cosine similarity based JL lemma can be improved by applying restricted isometry property (RIP). Let $\mathbf{A}$ be a matrix satisfying $(k,\delta)$-RIP, that is
@@
for all $k$-sparse $\mathbf{s}\in\mathbb{S}^{M-1}$, i.e., $\|\mathbf{s}\|_{0}\leq k$. The following result is established in [86].
@@
Let $\eta,\epsilon\in(0,1)$, $\{\mathbf{s}_{1},\ldots,\mathbf{s}_{K}\}\in\mathbb{S}^{M-1}$, and $\mathbf{A}\in\mathbb{R}^{m\times M}$ be $(k,\delta)$-RIP with $\delta\leq\epsilon/4$ and $k\geq 40\log\frac{4K}{\eta}$. Let $\bm{\sigma}$ a Rademacher sequence, i.e., uniformly distributed on $\{-1,1\}^{M}$. Then, with probability exceeding $1-\eta$,
$|\mathbf{s}_{i}^{T}\mathbf{s}_{j}-\mathbf{s}_{i}^{T}\mathbf{D}_{\bm{\sigma}}\mathbf{P}\mathbf{D}_{\bm{\sigma}}\mathbf{s}_{j}|\leq\epsilon,\quad\forall i,j\in\{1,\ldots,K\},$ (37)
@@
According to the theory of compressive sensing, the $m\times M$ partial Gaussian matrix can be used with
$m\geq\frac{C}{\epsilon^{2}}\log\frac{K}{\eta}\log M,$ (38)
@@
Recalling Definition 12, the distortion of semantic compression can be evaluated by the distance of two semantic vector spaces. Let $\mathsf{S}$ be the original semantic vector space $\mathbb{S}^{M-1}$ and $\mathsf{S}^{\prime}$ be $\mathbb{S}^{m-1}$ with $1\leq m<M$, the distortion of semantic compression can be written as
$d_{s}^{2}(\mathsf{S},\mathsf{S}^{\prime})=\min_{\pi\in\Pi(\mu,\mu^{\prime})}\int_{\mathbb{S}^{M-1}\times\mathbb{S}^{m-1}}\int_{\mathbb{S}^{M-1}\times\mathbb{S}^{m-1}}\left|\mathbf{s}^{T}\mathbf{s}^{\prime}-\mathbf{s}^{T}\mathbf{P}\mathbf{s}^{\prime}\right|^{2}d\pi(\mathbf{s},\mathbf{As})d\pi(\mathbf{s}^{\prime},\mathbf{As}^{\prime}),$ (39)
where $\mathbf{A}$ is a $m\times M$ projection matrix and $\mathbf{P}=\mathbf{A}^{T}\mathbf{A}$. The following theorem can be established by applying Lemma 1 or Theorem 3 directly.
@@ IV-D Semantic Embedding/Vectorization for Next-token Prediction
In practice, we typically select a proper dimension $m$ to directly perform the semantic embedding or vectorization. In the following, we will discuss information-theoretically optimal approach. It is natural to understand that the semantics of an utterance highly depend on the speaker’s intended goal, i.e., the downstream task in machine learning. Therefore, for a token sequence with length $n$, the semantic embedding is a mapping $f:\Omega^{n}\to(\mathbb{S}^{m-1})^{n}$, such that a loss functional $\mathcal{L}(f)$, defined by the downstream task, is minimized.
From the perspective of LLMs, the objective is to predict the next token based on the prompt and the parameterized memory. Therefore, $\mathcal{L}(f)$ should be designed to best facilitate of achieving this goal. Let $X_{1:n}$ be a token sequence, $S_{1:n}$ be the corresponding semantic vector representation of $X_{1:n}$. For the task of the next token prediction, $S_{i}$ should contain all the information in $X_{1:i}$ which is useful to predict $X_{i+1:n}$. From the perspective of information theory, the optimal semantic encoder for next token prediction should be the solution of the following problem:
$\max_{S_{i}=f(X_{1:i})}I(X_{i+1:n};S_{i}|S_{1:i-1}),\quad 1\leq i\leq n\in\mathbb{N}.$ (40)
The condition means $S_{i}$ only contains new information for predicting $X_{i+1:n}$ which is not contained in $S_{1:i-1}$.
@@
$I^{*}(X_{n:1}\to S_{1:n})\leq\sum_{i=1}^{n}\max_{S_{i}=f(X_{1:i})}I(X_{i+1:n};S_{i})\leq\sum_{i=1}^{n}\sum_{k=1}^{n-i}\max_{S_{i}=f(X_{1:i})}I(X_{i+k};S_{i}).$ (42)
Inspired by the idea of predictive coding in information theory [87, 88], the CPC is proposed for semantic embedding in [48], which is also adopt in OpenAI [49]. Let $Z_{1:n}$ be the latent representation of $X_{1:n}$ with $Z_{i}=g_{\text{ENC}}(X_{i})$, $S_{1:n}$ be the semantic vector obtained by CPC, which is defined as $S_{i}=g_{\text{AR}}(Z_{1:i-1})$. The training process of CPC is to solve the following optimization problem:
$\sum_{k=1}^{n-i}\max_{S_{i}=f(X_{1:i})}I(X_{i+k};S_{i}).$ (43)
@@ V-A TV-VAR based AR-LLMs
Let $\mathbf{s}_{i}$ with $i=1,\ldots,n$ and $\mathbf{u}_{i}$ with $i=n+1,\ldots,N$ be sample vectors of random variables $S_{i}$ and $U_{i}$, respectively. To simplify the notation, we let $\mathbf{u}_{i}=\mathbf{s}_{i}$ for $i=1,\ldots,n$. We then have the following definition.
@@
$\mathbf{u}_{i}=\arg\mathrm{softmax}\left(\frac{1}{T}\tilde{\mathbf{u}}_{1:M}^{T}\Psi\left(\sum_{j=1}^{i-1}\mathbf{A}_{ij}\mathbf{u}_{j}\right)\right),\quad i=n+1,\ldots,N,$ (44)
where $\mathbf{A}_{ij}$ is the coefficient matrix, $\tilde{\mathbf{u}}_{1:M}$ are all possible token vectors in $\mathcal{S}(\Omega)$, $\Psi(\cdot)$ is a function, and $T$ is the sampling temperature.
In contrast to the standard VAR model [50], $\mathbf{A}_{ij}$ is time-variant, which is very difficult to estimate in practice. According to Jaynes’ maximum entropy principle [89], the $\mathrm{softmax}$ is a probability assignment on discrete sample space that maximize the entropy with the constraint on the first order moment. Therefore, the obtained estimation of $\mathbf{u}_{i}$ is the one with the maximum uncertainty, i.e., the best achievable estimation in the worst case.
Inspired by statistical physics [90, 91], Hinton introduced the energy-based model (EBM) for neural networks, especially restricted Boltzmann machine [92, 93]. Similarly, the AR-LLM defined in Eq. (44) can also be interpreted from the EBM perspective. The probability distribution of $\mathbf{u}_{i}$ can be written as the following Boltzmann distribution:
$P(\mathbf{u}_{i})=\frac{1}{Z}e^{-\frac{1}{T}E(\mathbf{u}_{i})}$ (45)
where $T$ is the temperature in energy function, $Z$ is the partition function, defined as
$Z=\sum_{\mathbf{u}_{i}\in\mathcal{S}(\Omega)}e^{-\frac{1}{T}E(\mathbf{u}_{i})}.$ (46)
The energy function is defined as
$E(\mathbf{u}_{i})=-\left\langle\mathbf{u}_{i},\Psi\left(\sum_{j=1}^{i-1}\mathbf{A}_{ij}\mathbf{u}_{j}\right)\right\rangle.$ (47)
Eq. (44) can then be derived by finding $\mathbf{u}_{i}$ from $\mathcal{S}(\Omega)$ that maximizes the Boltzmann distribution, i.e., minimizes the energy function $E(\mathbf{u}_{i})$.